# Tactile Displays Driven by Projected Light

### v1

@@
Max Linnander1, Dustin Goetz1, Gregory Reardon1, Elliot Hawkes1, and Yon Visell1∗
1Department of Mechanical Engineering, Biological Engineering Program,
Media Arts & Technology Program,
@@
Tactile displays that lend tangible form to digital content could profoundly transform how we interact with computers, much like visual displays have driven successive revolutions in computing over the past 60 years. However, creating tactile displays with the actuation speeds, dynamic ranges, and resolutions that are required for perceptual fidelity has proved challenging. Here, we present a tactile display that directly converts projected light into visible tactile patterns using an energetically passive, photomechanical surface populated with arrays of millimeter-scale optotactile pixels. The pixels transduce incident light into mechanical displacements through rapid, light-stimulated thermal gas expansion, yielding displacements of up to 1 millimeter and response times of 2 to 100 milliseconds. Our use of projected light for power transmission and addressing enables these displays to be scaled in size and resolution at sustainable cost and complexity. We demonstrate devices with up to 1,511 independently addressable pixels. Perceptual studies confirm the capacity of the display to accurately reproduce tactile patterns in location, timing, frequency, and structure. This research establishes a foundation for practical, versatile high-resolution tactile displays driven by light.
@@ Introduction
Tactile displays give physical form to digital content, conveying information through tactile patterns that can represent data, user interfaces, or virtual objects. They could profoundly transform human-computer interaction, much like visual displays have driven successive revolutions in computing, from the first widespread computers of the 1970s to wearable and virtual reality systems of today  (?). Applications of tactile displays span many fields, including automotive interfaces that emulate physical controls with software reconfigurability  (?, ?, ?), educational tools such as animated haptic books  (?, ?), and user interfaces for virtual and augmented reality  (?, ?). However, creating multipurpose tactile displays with perceptual fidelity imposes stringent requirements on spatial resolution ($\lesssim$ 3 mm), response time ($<$50 ms), displacement ($>$ 0.2 mm), and force ($>$ 5 mN), with proportional output  (?, ?, ?, ?). Achieving these requirements in displays of useful size at sustainable costs will require significant advancements in actuation technologies, scalability, integration density, and fabrication methods.
@@
Despite this potential, the use of targeted light transmission for tactile actuation has been sparsely investigated. Prior studies have employed thermal mechanisms that convert absorbed optical energy into forces and displacements. Examples include solid thermoelastic expansion  (?), phase transitions in liquid crystal elastomers  (?, ?, ?), and liquid vaporization  (?). In phase change mechanisms, a substantial portion of input energy is consumed to overcome latent heat barriers, delaying actuation. Furthermore, thermal processes often yield slow response times – on the order of seconds or more – due to constraints imposed by heat transfer (Table S2). The characteristic timescale for heat transfer through an actuator of dimension $L$ is of order $\tau=L^{2}/\alpha$, where $\alpha$ is thermal diffusivity. For polymers and liquids, $\alpha$ is on the order of 10-7 m2/s. In contrast, gases possess thermal diffusivities that are two orders of magnitude higher, because diffusivity is inversely proportional to density. For this reason, gases may enable sufficiently fast refresh times to enable dynamic tactile displays driven by projected light.
@@
Building upon this insight, we present a tactile display that uses projected light to activate tactile patterns across a surface populated with small scale ($L$ = 3 mm) gas-filled pixels. Pixels are activated through the photostimulated heating and transient expansion of the encapsulated gas (Fig. 1). Light pulses with energies of 2.5 to 125 mJ are sufficient to activate localized displacements of up to 1 mm and forces of up to 55 mN, with response times of 2 to 100 ms. The actuation method, optical addressing technique, and simplicity of fabrication make this technique practical and scalable. We present implementations with up to 1,511 independently addressable pixels distributed over a $15\times 15$ cm2 area – several times more pixels than any previous tactile display approximating the requirements cited above (Fig. 1H, Table S1). We show that these light-driven tactile displays are capable of representing a wide variety of spatiotemporal tactile patterns that can be accurately perceived.
@@
Figure 1: Fig. 1. Tactile display driven by projected light. A) Projected light excites tactile patterns, consisting of mechanical signals produced by an array of optotactile pixels. B) Light is converted to forces and displacements via an assembly of simple patterned layers, easing manufacturing and scalability. From top: Elastic membrane (EcoFlex 00-10), cavity walls (polysiloxane, PS), thin photoabsorber (PGS, thickness: 17 $\mu$m), cavity walls (PS), optically transparent layer (acrylic). C) Cross-section view of pixel operation. An incident light pulse is converted into heat by the photoabsorber. Heat is transferred to air in the cavity, increasing gas temperature $T_{gas}$ and pressure $P$. Gas expansion drives the deflection of the elastic membrane. D) Finite element analysis of pixel thermal and mechanical response. E) Flexible tactile display with 357 pixels. Scale bar: 1 cm. F) Photostimulation yields both tactile and visual feedback, due to the transmission of scattered light through the membrane. Extended exposure image of 1,511 pixel display. Scale bar: 1 cm. G) The same technique can be used to realize displays that vary in dimensions and pixel count. From upper left: 1, 8, 357, and 1511 pixels. Scale bar: 1 cm. H) Comparison with prior tactile displays. The projected light tactile display in this work yields fast response rates, scales to pixel counts exceeding prior displays, and is uniquely capable of both visual and tactile display. Driving mechanisms: L = light, EM = electromagnetic, PN = pneumatic, ES = electrostatic, T = thermal, PZ = piezoelectric, C = combustion, and EO = electroosmosis. (See also Fig. S16, Table S1).
@@
The display functions as an interface that translates projected light into tactile mechanical signals, via an array of optotactile pixels (Fig. 1A). Each pixel operates on a thermodynamic principle, transducing heat from an absorbed light pulse into mechanical energy. Light entering the pixel via an optical window is converted to heat by a thin photoabsorber, which rapidly increases in temperature (Fig. 1C). Heat is transferred to gas within the cavity, causing a pressure increase dictated by the ideal gas law $PV=nRT_{gas}$. Gas expansion drives the deflection of an elastic membrane sealing the pixels, yielding localized mechanical forces $F(t)$ and displacements $d(t)$ that can be felt by touching the membrane (Fig. 1D).
@@
This general strategy may be used to realize a large variety of displays. The following design choices exemplify one way it can be implemented. A multi-layered architecture was used to facilitate ease of manufacturing and scalability (Fig. 1B). Transparent acrylic forms a base layer, functioning as the optical window (Fig. S2). The base is adhered to two polysiloxane (PS) layers with arrays of cylindrical cavities (3 mm diameter, 1 mm height) (Fig. 1C, S4-5). Between these layers is a thin pyrolytic graphite sheet (PGS, thickness: 17 $\mu$m) that is patterned to form an array of photoabsorbers, which are suspended within each cavity by narrow bridge features (Fig. 1C-D,S4). PGS was selected due to its high thermal conductivity, optical absorptance ($\varepsilon$=65%; Fig. S2), and thermal stability [^2] PGS is stable up to 3000 C; Sustained temperatures over 400 C may favor the use of a more inert gas, such as nitrogen or helium, to inhibit oxidization.. Capping this assembly is the tactile surface – an elastomer membrane that seals the air-filled cavities. The display can be made flexible through the use of a thinner, or more elastic, base layer (Fig. 1E). The entire layered assembly is energetically passive.
@@
The surface is stimulated by brief (1 to 50 ms) low-energy light pulses (2.5 to 125 mJ) delivered from a compact diode laser ($\lambda=450$ nm; Fig. 1C,S6). Scanning is facilitated by commodity x-y galvanometer mirrors, yielding spatiotemporal tactile patterns that can be felt, and also seen, due to the transmission of scattered light through the membrane (Fig. 1F). The mechanism of actuation, optical addressing technique, simplicity of fabrication, and use of widely available materials render this technology practical and scalable. To demonstrate the ease of manufacturing and scalability, we fabricated devices with 1, 8, 357, and 1,511 addressable pixels (Fig. 1G). The pixel dimensions (diameter 3 mm, height 1 mm) were chosen based on theoretical analyses and finite-element numerical studies (Text S1, S2). As demonstrated below, our technique achieves forces, displacements, resolutions, and response rates meeting perceptual requirements for tactile display, at display scales suitable for diverse applications (Fig. 1H, Fig. S16, and Table S1).
@@
Figure 2: Fig. 2. Thermal-mechanical response characterization. A,B) Absorber surface temperature and membrane deflection increased monotonically during illumination with power $P_{L}=2.5$ W over 50 ms. (Scale bar = 0.5 mm). C) Reducing the thermal bridge width from $w=$ 0.75 mm to 0.2 mm yielded higher temperatures and larger displacements. Scale bar: 0.5 mm. D) Experimental absorber temperature $T(t)$ and membrane displacement $d(t)$ vs. time $t$ for $w=$ 0.2, 0.25, 0.4, 0.55, and 0.75 mm. As $w$ decreased, more heat was retained in the absorber, yielding higher temperatures and larger displacements. For $w=0.2$ mm, at $t=$ 50 ms we observed $T=527^{\circ}$ C and $d=0.97$ mm. Grey dashed line: Numerical results for $w=0.25$ mm; see Fig. S1. E) The thermal resistance $R$ of the bridge increased as width $w$ decreased. The data agreed with the theoretically predicted relationship $R=a/w,$ with $a=$ 73.3 mm K/W, $r^{2}=0.98$. F) Peak membrane displacement $d$ and absorber thermal relaxation time $\tau$ increased with thermal bridge resistance $R$, while the relaxation rate $1/\tau$ decreased, indicating that tuning $w$ involves a tradeoff between displacement and relaxation rate.
@@
We experimentally characterized the thermal and mechanical response of individual pixels under pulsed photostimulation at constant power $P_{L}$ for pulse length $t_{p}=50$ ms. We observed the mean temperature $T(t)$ in the photoabsorber to rise monotonically during exposure (Fig. 2A-D), driving expansion of air within the pixel cavity and deflection of the elastic membrane (Movie S1). Temperature and displacement both reached their maximum values near the time at which the optical pulse ceased, indicating a nearly static relationship between the absorber and air temperatures during heating (Fig. 2D, Fig. S8). Following the pulse, temperature and displacement relaxed toward baseline conditions.
@@
We studied the effect of photoabsorber geometry by varying the width, $w$, of the bridges suspending it in the cavity, from $w=$ 0.2 to 0.75 mm (Fig. 2A-D). Larger absorber temperatures, $T(t),$ and displacements, $d(t),$ are obtained from pixels with smaller widths, $w$, indicating that bridge width has a dominant effect on heat transfer. We confirmed this using theoretical and numerical analyses of heat transfer (Text S1-2). The highest absorber temperatures and displacements were obtained at the smallest width, $w=0.2$ mm. Temperature increased by $507$ C and displacement reached a maximum of $0.97$ mm after 50 ms of photostimulation at 2.5 W. The thermal and mechanical response of the pixels can thus be tuned by varying the bridge width, $w$.
@@
We analyzed the dynamic thermal response using the experimental data in tandem with a standard description of heat transfer (Text S1). An effective description of heat transfer indicates that the time scale $\tau$ governing the absorber temperature $T(t)$ response to photostimulation is given by $\tau=RC,$ where $C$ is the absorber heat capacity and $R$ is the effective thermal resistance of the absorber. We determined $R$ for each width $w$ by fitting the experimental data at the onset of photostimulation (Table S3). The effective thermal resistance was inversely proportional to bridge width, in agreement with the theoretically expected relationship $R=L(k_{xy}hw)^{-1}$ where $k_{xy}$ is the in-plane thermal conductivity of the PGS absorber, $h$ is thickness, and $L$ is the characteristic length of heat transfer (Fig. 2E, Text S1). $R$ thus provides an useful parameter for examining the thermal response of photoabsorbers with different geometries. Varying $R$ yielded thermal response time scale values in the range 10 ms $\leq\tau\leq$ 100 ms. Pixels with higher values of $R$ yielded greater displacements, $d$, and larger values of $\tau$ (Fig. 2F). These competing effects illustrate a trade off between response amplitude and speed that we examined in further experiments.
@@
Achieving rapid pixel response rates is essential for meeting the perceptual requirements of tactile displays. Refresh intervals shorter than 50 ms are necessary to ensure perceptual continuity of tactile patterns across actuation cycles  (?, ?), similar to frame rate requirements for persistence of vision in visual displays. Rapid response rates also enhance the bandwidth of information conveyance and increase display responsiveness. Our display exhibits rapid actuation cycles, from 2 to 100 ms, that contrast with slower response times, 1 or more seconds, that are often obtained with thermally mediated actuation (Fig. S16). This performance is enabled by the low thermal mass of the photoabsorber ($100$ $\mu$J/K) and encapsulated air ($10$ $\mu$J/K), by the small quantity of heat energy delivered in each pulse (2.5 to 125 mJ), and by favorable heat transfer characteristics, including the surface area to volume ratio, $A/V\propto 1/L$, where $L=3$ mm is the characteristic pixel dimension.
@@
However, achieving rapid actuation cycles is balanced by the competing need for sufficient dynamic range in force and displacement. This trade-off is parametrized by the effective thermal resistance, $R$, of the absorber. To evaluate this balance in our system, we experimentally characterized the membrane force and displacement responses across a range of optical stimuli (Fig. 3). We observed the peak output force to increase linearly with increasing optical power $P_{L}$ (Fig. 3B). At constant power, force and displacement increased monotonically with pulse length $t_{p}$ (Fig. 3B, Fig. S7). Photostimulation of higher-resistance pixels ($R=382$ K/W) yielded greater peak forces ($F=55$ mN) after $t_{p}=50$ ms, while lower resistance pixels ($R=145$ K/W) yielded smaller peak forces ($F=31$ mN) (Fig. 3B; Fig. S9).
@@
Using these measurements, we computed the energetic efficiency, $\eta,$ of actuation as the ratio of stroke power to absorbed optical power, obtaining $\eta=0.03\%$ (Text S3, Fig. S11). This efficiency is within the range observed for many thermal actuation methods (Table S1). It reflects several losses. More than 90% is due to photoabsorber heat transfer to the walls via the bridges (Text S3). This thermal conduit speeds the pixel response time, but decreases efficiency. Adjusting $R$ allows us to tune performance, trading off efficiency and speed. Heat in the gas is converted to output work with about 14% efficiency. This loss is primarily due to entropy generation  (?), a measure of irreversibility, which is magnified at small scales. In our device, the pixel volume is $V=7.8$ $\mu$L. A theoretical scaling analysis indicates that when the characteristic dimension $L$ of a thermally actuated device is sufficiently small, the relative loss in efficiency $\eta$ scales as $\Delta\eta/\eta_{r}\propto L^{-\alpha}$, where $\eta_{r}$ is the Carnot efficiency, and $\alpha=1$ or $2$ reflects the thermodynamic design (Text S4). Similar size-efficiency effects have been experimentally observed across many thermodynamic systems  (?, ?).
@@
Figure 3: Fig. 3. Cyclic photostimulated actuation. A) Isometric force and unloaded displacement measurements. B) Individual pixels produced peak forces of $F=$ 55 mN under photostimulation at power $P_{L}=2.5$ W and pulse duration $t_{p}=50$ ms (pulse energy 125 mJ). Force increased linearly with power $P_{L}$ and increased with pulse duration $t_{p}$. C) Successive displacement responses to pulsed excitation exhibit minimal variation for pulse gap intervals $t_{g}\gg\tau$ (Fig. S7), where $\tau$ is the thermal relaxation time; Here, $\tau\approx 23$ ms and $R=145$ K/W. At shorter intervals $t_{g},$ residual heat yields a DC offset $\delta_{o},$ reducing peak-to-peak displacement $\delta_{pp}$ relative to initial displacement $\delta_{1}$. D) Steady-state amplitude $\delta_{pp}/\delta_{1}$ (blue) increased with increasing pulse gap $t_{g}$. Blue line: $\delta_{pp}/\delta_{1}=1-\exp{(t_{g}/\tau)}$ ($r^{2}=0.95$). Normalized offset $\delta_{o}/\delta_{1}$ (red) followed an opposite trend ($r^{2}=0.81$). (Uncertainty bars = 1 SD, from $n=5$ oscillations). E) Displacements $\delta_{pp}$ for response rates up to $f=$ 500 Hz. Peak-to-peak displacement decreased as $f=1/(t_{p}+t_{g})$ increased ($t_{p}/t_{g}=1$). Inset: displacement $d(t)$ in steady-state for $f=20$ Hz and $200$ Hz. F) In sequential scanning of pixels, the maximum pixel display rate $N$ is dictated by the required displacement. At fixed $P_{L}$, $N$ decreases as displacement increases. For $P_{L}=2.5$ W, $N=$ 217 pixels/s is achieved at 50 $\mu$m, and $N=26.5$ pixels/s is achieved at 400 $\mu$m.

### v3

@@
Max Linnander1, Dustin Goetz1, Gregory Reardon3,
Vijay Kumar1, Elliot Hawkes1, and Yon Visell1,2,3,4∗
1Department of Mechanical Engineering, 2Department of Bioengineering,
3Media Arts & Technology Program, 4Department of Electrical and Computer Engineering,
@@
Tactile displays that lend tangible form to digital content could transform computing interactions. However, achieving the resolution, speed, and dynamic range needed for perceptual fidelity remains challenging. We present a tactile display that directly converts projected light into visible tactile patterns via a photomechanical surface populated with millimeter-scale optotactile pixels. The pixels transduce incident light into mechanical displacements through photostimulated thermal gas expansion, yielding millimeter scale displacements with response times of 2 to 100 milliseconds. Employing projected light for power transmission and addressing renders these displays highly scalable. We demonstrate optically driven displays with up to 1,511 addressable pixels – several times more pixels than any prior tactile display attaining comparable performance. Perceptual studies confirm that these displays can reproduce diverse spatiotemporal tactile patterns with high fidelity. This research establishes a foundation for practical, versatile high-resolution tactile displays driven by light.
One-Sentence Summary
Photomechanical surfaces directly convert projected light into visible and tactile patterns.
@@ Introduction
Tactile displays give physical form to digital content, conveying information through tactile patterns that can represent data, user interfaces, or virtual objects. They could profoundly transform human-computer interaction, paralleling the role of visual displays in driving successive revolutions in computing, from the first widespread computers of the 1970s, to the mobile, wearable, and virtual reality systems of today  (?). Applications of tactile displays span many fields, including automotive interfaces that emulate physical controls with software reconfigurability  (?, ?, ?), educational tools such as animated haptic books  (?, ?), and user interfaces for virtual and augmented reality  (?, ?). However, creating multipurpose tactile displays with perceptual fidelity imposes stringent requirements on spatial resolution ($\lesssim$ 3 mm), response time ($<$50 ms), displacement ($>$ 0.2 mm), and force ($>$ 5 mN), with proportional output  (?, ?, ?, ?). Achieving these requirements in displays of useful size at sustainable costs will require significant advancements in actuation technologies, scalability, integration density, and fabrication methods.
@@
Despite its potential, targeted light transmission for tactile actuation has been sparsely investigated. Prior studies have employed thermal mechanisms for photomechanical transduction, by converting absorbed light into heat, which is then transformed into tactile actuation. Examples include solid thermoelastic expansion  (?), phase transitions in liquid crystal elastomers  (?, ?), and liquid vaporization  (?). Due in part to constraints arising from heat transfer, these systems have often exhibited slow response times of seconds or longer (Table S2). Employing a captive gas as the working medium can enable tactile displays to achieve the rapid response times that are required for perceptual fidelity. Gases exhibit high thermal diffusivity, $\alpha$, due to their low density. High diffusivity promotes rapid heat transfer, because the heat transfer timescale is given by $\tau=L^{2}/\alpha$, where $L$ is a characteristic length. The low density of gases also results in large temperature increases $\Delta T=Q/(c\rho V)$, where $Q$ is heat input, $V$ is volume, and $c$ is specific heat capacity. Larger values of $\Delta T$ result in faster thermal responses because heat is transferred out of the actuator at a rate described by Newton’s law of cooling, $\dot{Q}=hA\Delta T$, where $h$ is a heat transfer coefficient and $A$ is effective surface area.
Directly heating a captive gas with light is challenging, as gases are generally transparent. However, efficient heat transfer to a gas can be achieved by integrating a low thermal mass photoabsorber within a gas-filled cavity. Such a configuration may enable rapid actuation cycles suitable for tactile displays driven by projected light. This principle has historical parallels in experiments by Bell, Tyndall, and Röntgen in the 1880s, where intermittent illumination of absorbing materials in gas-filled cavities was observed to yield audible sound  (?, ?, ?). Similar principles have been exploited in prior research on remote power transmission for aerospace  (?), thermodynamic actuation  (?), and optically driven micromechanical valves  (?, ?, ?), among other systems.
@@
Building upon these insights, we present a tactile display that uses projected light to activate tactile patterns across a surface populated with small optotactile pixels ($L$ = 3 mm). Each pixel contains a thin film photoabsorber suspended in a captive gas. Photostimulation of a pixel drives rapid heating, inducing transient gas expansion (Fig. 1). Light pulses with energies of 2.5 to 125 mJ are sufficient to activate localized displacements of up to 1 mm and forces of up to 55 mN, with response times of 2 to 100 ms. The actuation method, optical addressing technique, and simplicity of fabrication make this technique practical and scalable. We present implementations with up to 1,511 independently addressable pixels distributed over a $15\times 15$ cm2 area – several times more pixels than any previous tactile display approximating the requirements cited above (Fig. 1H, Table S1). We show that these light-driven tactile displays are capable of representing a wide variety of spatiotemporal tactile patterns that can be accurately perceived.
@@
Figure 1: Fig. 1. Tactile display driven by projected light. (A) The display directly converts projected light into visible, tactile patterns via an array of optotactile pixels. (B) Light is converted to forces and displacements via an assembly of patterned layers, supporting manufacturing and scalability. From top: Elastic membrane (EcoFlex 00-10), cavity walls (polysiloxane, PS), thin photoabsorber (PGS, thickness: 17 $\mu$m), cavity walls (PS), optically transparent layer (acrylic). (C) Pixel operation (section view). An incident light pulse is converted into heat by the photoabsorber. Heat is transferred to air in the cavity, raising gas temperature $T_{gas}$ and pressure $P$. Gas expansion drives the deflection of the elastic membrane. (D) Finite element analysis of pixel thermo-mechanical response. (E) Flexible tactile display, 357 pixels. Scale bar: 1 cm. (F) Photostimulation elicits both tactile and optical feedback, due to the transmission of scattered light through the membrane. Extended exposure image of 1,511 pixel display. Scale bar: 1 cm. (G) This technique can be used to realize displays that vary in dimensions and pixel count. From upper left: 1, 8, 357, and 1,511 pixels. Scale bar: 1 cm. (H) Comparison with prior tactile displays (Fig. S15, Table S1). Our display enables rapid response rates, scales to pixel counts surpassing prior work, and is uniquely capable of both visual and tactile display. Driving mechanisms: L = light, EM = electromagnetic, PN = pneumatic, ES = electrostatic, T = thermal, PZ = piezoelectric, C = combustion, and EO = electroosmosis.
@@
The display reproduces visible, tactile patterns by translating projected light into mechanical signals, using an array of optically addressed pixels (Fig. 1A,B). Each pixel operates on a thermodynamic principle, transducing optical energy into heat, and heat into work. Light entering the pixel is converted to heat by a thin photoabsorber suspended in a captive gas (Fig. 1C). The absorber rapidly increases in temperature, transferring heat to the gas, which undergoes a pressure increase dictated by the ideal gas law $PV=nRT_{gas}$. Gas expansion drives the deflection of an elastic membrane sealing the pixels, yielding localized mechanical forces $F(t)$ and displacements $z(t)$ that can be felt by touching the membrane (Fig. 1D).
@@
This general strategy may be used to realize a large variety of displays. The following design choices exemplify one way it can be implemented. A multi-layered architecture was used to facilitate ease of manufacturing and scalability (Fig. 1B). Transparent acrylic forms a base layer, functioning as the optical window (Fig. S2). An array of cylindrical pixel cavities is formed from two polysiloxane (PS) layers that are patterned via laser cutting (Fig. 1C, S4-5). The pixel dimensions (diameter 3 mm, height 1 mm) were selected based on perceptual requirements, theoretical analysis, and numerical simulations (Text S1, S2). Between the pixel cavity layers is a pyrolytic graphite sheet (PGS, 17 $\mu$m) that is patterned to form an array of photoabsorbers suspended by narrow bridge features (Fig. 1C-D, S4). PGS was selected due to its high thermal conductivity, broad spectrum photoabsorption (Fig. S2), and thermal stability [^2] PGS is stable up to 3000 C; Sustained temperatures over 400 C may favor the use of a more inert gas, such as nitrogen or helium, to inhibit oxidization.. Capping this assembly is the tactile surface – an elastomer membrane that seals the air-filled cavities. The display can be made flexible through the use of a thinner or more elastic transparent base layer (Fig. 1E).
@@
The pixels are stimulated by brief low-energy light pulses (2.5 to 125 mJ) delivered from a compact diode laser ($\lambda=450$ nm; Fig. 1C, S6). Optical addressing is facilitated by commodity xy scanning galvanometer mirrors, yielding spatiotemporal tactile patterns that can be felt and also seen. The patterns are made visible through the transmission of light scattered through the membrane at each pixel (Fig. 1F). The mechanism of actuation, optical powering and addressing technique, simplicity of fabrication, and use of widely available materials render this technology practical and scalable (Table S5, S6). To demonstrate the ease of manufacturing and scalability, we fabricated devices with 1, 16, 357, and 1,511 addressable pixels (Fig. 1G), using the same fabrication methods for each. Our results indicate that the devices achieve forces, displacements, resolutions, and response rates meeting perceptual requirements for tactile display, at display scales suitable for diverse applications (Fig. 1H, Fig. S15, and Table S1).
@@
We experimentally characterized the thermal and mechanical response of individual pixels driven by discrete light pulses with duration $t_{p}=50$ ms and power $P_{L}$. We observed the mean temperature $T(t)$ in the photoabsorber to rise monotonically during exposure (Fig. 2A-D), driving expansion of air within the pixel cavity and deflection of the elastic membrane (Movie S1). Temperature and displacement both reached their maximum values near the time at which the optical pulse ceased, indicating a nearly static relationship between the absorber and air temperatures during heating (Fig. 2D, Fig. S8). Following the pulse, temperature and displacement relaxed toward baseline conditions.
Figure 2: Fig. 2. Thermal-mechanical response under photostimulation. (A,B) Absorber surface temperature and membrane deflection increased monotonically during illumination with power $P_{L}=2.5$ W over 50 ms. (Scale bar = 0.5 mm). (C) Reducing the thermal bridge width from $w=$ 0.75 mm to 0.2 mm yielded higher temperatures and larger displacements. Scale bar: 0.5 mm. (D) Experimental absorber temperature $T(t)$ and membrane displacement $z(t)$ vs. time $t$ for $w=$ 0.2, 0.25, 0.4, 0.55, and 0.75 mm. As $w$ decreased, more heat was retained in the absorber, yielding higher temperatures and larger displacements. For $w=0.2$ mm, at $t=$ 50 ms we observed $T=527$ C and $d=0.97$ mm. Grey dashed line: Numerical results for $w=0.25$ mm; see Fig. S1. (E) The thermal resistance $R$ of the bridge increased as width $w$ decreased. The data agreed with the theoretically predicted relationship $R=a/w,$ with $a=$ 73.3 mm K/W, $r^{2}=0.98$. (F) Peak membrane displacement and absorber thermal relaxation time $\tau$ increased with thermal bridge resistance $R$, while the relaxation rate $1/\tau$ decreased, indicating that tuning $w$ involves a tradeoff between displacement and relaxation rate.
@@
The thermal and mechanical response characteristics of the pixel can be tuned by adjusting the photoabsorber geometry. We studied the effect of varying the width, $w$, of the bridges suspending the photoabsorber in the cavity, from $w=$ 0.2 to 0.75 mm (Fig. 2A-D, Movie S2). Larger absorber temperatures, $T(t),$ and displacements, $z(t)$, are obtained from pixels with smaller widths, $w$, indicating that bridge width has a dominant effect on heat transfer out of the pixel. We confirmed this using theoretical and numerical analyses of heat transfer (Text S1-2). The highest absorber temperatures and displacements were obtained at the smallest width, $w=0.2$ mm. Temperature increased by $507$ C and displacement reached a maximum of $0.97$ mm after 50 ms of photostimulation at 2.5 W (Movie S3).
@@
A thermal resistance parameter, $R$, encodes the influence of photoabsorber bridge width, $w,$ on heat transfer, and can be adjusted to tune pixel response time or displacement. We identified $R$ by analyzing the experimental data in tandem with an effective description of heat transfer (Text S1). The time scale $\tau$ governing the absorber temperature $T(t)$ response to photostimulation is given by $\tau=RC,$ where $C$ is the absorber heat capacity. Values for $R$ and $C$ were obtained by numerical fit to experimental data (Table S3). $R$ is inversely proportional to $w$, in agreement with the theoretically expected relationship $R=L(k_{xy}hw)^{-1}$ where $k_{xy}$ is the in-plane thermal conductivity of the PGS absorber, $h$ is thickness, and $L$ is the characteristic length of heat transfer (Fig. 2E, Text S1). By adjusting $R,$ we can tune the thermal response time in the range 10 ms $\leq\tau\leq$ 100 ms. Selecting larger values of $R$ yields larger peak displacements, and longer response times, $\tau$ (Fig. 2F). These competing effects illustrate a trade off between response amplitude and speed that we examined in further experiments.
@@
Achieving rapid pixel responses is essential for meeting the perceptual requirements for cyclically refreshed tactile display. Refresh intervals shorter than 50 ms are required in order to ensure perceptual continuity of tactile patterns across actuation cycles  (?, ?), similar to frame rate requirements for persistence of vision in visual displays. Rapid response rates also enhance the bandwidth of information conveyance and increase display responsiveness. Our display exhibits rapid actuation cycles, from 2 to 100 ms, that contrast with slower response times, 1 or more seconds, that are often obtained with thermally mediated actuation (Fig. S15). This performance is enabled by the high thermal diffusivity of the captive gas, and low thermal mass of the photoabsorber ($100$ $\mu$J/K) and gas ($10$ $\mu$J/K), by the small quantity of heat energy delivered in each pulse (2.5 to 125 mJ), and by favorable heat transfer characteristics, including the surface area to volume ratio, $A/V\propto 1/L$, where $L=3$ mm is the characteristic pixel dimension.
@@
However, achieving rapid actuation cycles is balanced by the competing need for sufficient dynamic range in force and displacement. This trade-off is parametrized by the effective thermal resistance of the absorber, $R$. To evaluate this balance in our system, we experimentally characterized the membrane force and displacement responses across a range of optical stimuli (Fig. 3). We observed the peak output force to increase linearly with increasing optical power $P_{L}$ (Fig. 3B). At constant power, force and displacement increased monotonically with pulse duration $t_{p}$ (Fig. 3B, Fig. S7). With pulse duration held constant, we observed the largest forces ($F=55$ mN for $t_{p}=50$ ms) at the highest resistance ($R=382$ K/W), and smaller forces ($F=31$ mN) at lower resistance ($R=145$ K/W; Fig. 3B, S9).
@@
Using these measurements, we computed the stroke efficiency, $\eta_{s},$ of actuation as the ratio of stroke power to absorbed optical power, obtaining $\eta=0.03\%$ (Text S3, Fig. S11). This efficiency is within the range observed for many thermal actuation methods (Table S1). It reflects several losses. More than 90% is due to photoabsorber heat transfer to the walls via the bridges (Text S3). This thermal conduit can facilitate rapid pixel responses, at the expense of efficiency and dynamic range. Adjusting $R$ allows us navigate this balance. The thermodynamic conversion of heat in the gas to work is about 14% efficient. The loss represents entropy generation  (?), a measure of irreversibility, and is magnified at small scales. In our device, the pixel volume is $V=7.8$ $\mu$L. A general scaling analysis indicates that when the characteristic dimension $L$ of a thermally actuated device is sufficiently small, the loss in efficiency scales as $\Delta\eta/\eta_{r}\propto L^{-\alpha}$, where $\eta_{r}$ is the Carnot efficiency, and $\alpha=1$ or $2$ depending on the heat input configuration (Text S4). Similar size-efficiency effects have been experimentally observed across many thermodynamic systems  (?, ?).
@@
Figure 3: Fig. 3. Cyclic photostimulated actuation. (A) Isometric force and unloaded displacement measurements. (B) Individual pixels produced peak forces of $F=$ 55 mN under photostimulation at power $P_{L}=2.5$ W and pulse duration $t_{p}=50$ ms (pulse energy 125 mJ). Force increased linearly with power $P_{L}$ and increased with pulse duration $t_{p}$. (C) Successive displacement responses to pulsed excitation exhibit minimal variation for pulse gap intervals $t_{g}\gg\tau$ (Fig. S7), where $\tau$ is the thermal relaxation time; Here, $\tau\approx 23$ ms and $R=145$ K/W. At shorter gap intervals, $t_{g},$ heat accumulates, yielding a slowly varying component, $d(t),$ and a reduction in the peak-to-peak oscillation amplitude, $\delta_{pp},$ relative to the first pulse amplitude, $d_{1}$. (D) As $t_{g}$ increases, oscillation amplitude, $\delta_{pp},$ (blue) increases; Regression fit: $\delta_{pp}/d_{1}=1-\exp{(t_{g}/\tau)}$ ($r^{2}=0.95$). The magnitude, $d$, of the slow component (red) followed an opposite trend, $d/d_{1}=\tau/t_{g}$ ($r^{2}=0.81$). (Uncertainty bars: 1 SD, $n=5$ trials). (E) Displacements $\delta_{pp}$ for pulse frequencies $f=$ 5 to 500 Hz. Peak-to-peak displacement decreased as $f$ increased, with $t_{p}/t_{g}=1$. Insets: oscillating component, $\delta(t),$ for $f=20$ Hz (yellow) and $200$ Hz (blue). (F) In sequential scanning, the maximum pixel display rate, $N,$ is dictated by the required displacement. At fixed $P_{L}$, $N$ decreases as displacement increases. For $P_{L}=2.5$ W, $N=$ 217 pixels/s is achieved at 50 $\mu$m, and $N=26.5$ pixels/s is achieved at 400 $\mu$m.